If a+b+c=3,a^2+b^2+c^2=13,a^3+b^3+c^3=27Then find the values of a,b,c
 

Grade:8

1 Answers

Sirajuddin
5 years ago
45:
a+b+c=3
(a+b+c)2-2(ab+bc +ac)=13
32-2(ab+bc+ac)=13
ab +bc+ac=-2
Now, a3+b3+c3=27
(a+b+c)(a2+b2+c-ab-bc-ac)